OpenStudy (anonymous):

please please please help,What is the equation of the parabola, in vertex form, with vertex at (-2,-4) and directrix x = -6?

OpenStudy (anonymous):

we can do this, just give me a second to make sure my answer is correct

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

good thing i checked, it wasn't let me start again

OpenStudy (anonymous):

alright no problem thank you for helping

OpenStudy (anonymous):

ok i made mistake in writing it but now i have it

OpenStudy (anonymous):

first of all since \(x=-6\) is a vertical line, and the vertex is to the right of the directrix you know it looks like this 

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

crappy picture but you get the idea we need to know that because that tells you that the \(y\) term is squared

OpenStudy (anonymous):

therefore the standard form of this will be \[(y+k)^2=4p(x-h)\] where the vertex is \((h,k)\) so we know right away that it is \[(y+4)^2=4p(x+2)\] now all we need is \(p\)

OpenStudy (anonymous):

typo there, standard form is \[(y-k)^2=4p(x-h)\]

OpenStudy (anonymous):

i really appreciate the help and explanation i suck at this

OpenStudy (anonymous):

is it okay so far? vertex is \((-2,-4)\) so we know it must be \[(y+4)^2=4p(x+2)\]

OpenStudy (anonymous):

now to find \(p\) the vertex is \((-2,-4)\) and the directix is \(x=-6\) the distance between \(-6\) and \(-2\) is \(4\) i.e. the directrix is 4 units to the left of the vertex this tell you \(p=4\) and so \(4p=16\)

OpenStudy (anonymous):

therefore your parabola is \[(y+4)^2=16(x+2)\]

OpenStudy (anonymous):

that helped so much thank you :)

OpenStudy (anonymous):

you can check (like i did) that this is correct here http://www.wolframalpha.com/input/?i=parabola+%28y%2B4%29^2%3D16%28x%2B2%29

OpenStudy (anonymous):

yw